An acquirer with a P/E ratio of 15 and earnings of $1.12 seeks to take over another target firm with value $16.32 and P/E ratio 14. What is the new merged firm's P/E?

(Assume a 1 to 1 share equivalence.)
Answer #1

Assume that there are 2 companies A and B

A:-

P/E ratio = 15 Earning= 1.12

B:-

P/E ratio= 14 and Value= 16.32

The Earning of co. B would be 16.32/14= 1.16

Both A and B is accuired at 1:1 ratio.

Therefore the merged P/E would be

16.32/(1.16+1.12)/2

14.31